[Hallc_running] [New Logentry] RC daily update
mnycz at jlab.org
mnycz at jlab.org
Tue Aug 18 11:30:01 EDT 2020
During SWING and OWL shifts we have continued with production on the polarized He3 cell.
RF recovery began at the start of DAY shift. While beam was down, we dropped to controlled access in order to perform opportunistic work in the Hall. This included
-Work on the solid Cherenkov detector
- Check the voltage in the hall under full load conditions
- Check drip bags on the top of the SHMS
Also, during the RF recovery, a temperature study at 90 degrees was performed -- https://logbooks.jlab.org/entry/3829879
Once the temperature study was completed, the holding field was rotated to 270 degrees -- https://logbooks.jlab.org/entry/3829873
When changing the QWP values, the TO noticed it wasn't responding -- https://logbooks.jlab.org/entry/3829861
It was rebooted following Ethan's procedure -- https://logbooks.jlab.org/entry/3829876
As we return from controlled access to beam permit, we will wait for the polarization to increase.
We will follow the short beam checkout procedure once MCC is ready to return sending beam to the Hall -- https://hallcweb.jlab.org/wiki/index.php/Beam_and_Raster_Checkout_for_A1n/d2n_(Short_Procedure)
Depending on the time MCC recovers beam to the Hall, we will likely take reference cell runs while the polarization increases.
Finally, as Brad mentioned in his email, please keep in mind the COVID-19 restrictions in place in the counting house to keep ensure everyone stays healthy.
We will meet today at 4:15pm via Bluejeans 861439920 -- https://hallcweb.jlab.org/wiki/index.php/RC_Update_and_Run_Plan_for_Aug_18,_2020
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3829919
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running